# Cyclotomic generating functions

### v2

@@ 3.5. CGF’s with a fixed denominator
Lemma 3.14.
Fix a multiset $B=\{b_{1},\ldots,b_{m}\}$ of positive integers. Suppose $\{a_{1},\ldots,a_{m}\}$ and $\{a_{1}^{\prime},\ldots,a_{m}^{\prime}\}$ are multisets of positive integers such that both $\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}$ and $\prod_{k=1}^{m}[a_{k}^{\prime}]_{q}/[b_{k}]_{q}\in\Phi^{+}$ are cyclotomic generating functions. Then there exists a multiset $\{h_{1},\ldots,h_{m}\}$ of positive integers such that for all $1\leq i\leq m$,
[…]
@@
By long division, one may observe $[xy]_{q}/[y]_{q}=[x]_{q^{y}}\in\Phi^{+}$. So, we may replace any $[a_{k}]_{q}$ in the numerator of a cyclotomic generating function with $[\ell a_{k}]_{q}$ for any positive integer $\ell$ and get another cyclotomic generating function. Hence, defining $h_{k}=a_{k}a_{k}^{\prime}$ for $1\leq k\leq m$ has the desired property. ∎
Corollary 3.15.
@@
4. Asymptotic considerations
Given a sequence of cyclotomic generating functions and their corresponding random variables, we can put their distributions in one common frame of reference by standardizing them to have mean 0 and standard deviation 1. Under what conditions does such a sequence converge? Can we classify all possible standardized limiting distributions of random variables coming from CGF’s? Both problems have been completely solved for certain families of CGF’s coming from $q$-hook length formulas, but the complete classification is not know, see [BS22, Open Problem 1.19]. In several “generic” regimes, they are asymptotically normal [BKS20a, Thm. 1.7], [BS22, Thm. 1.13]. In mildly degenerate regimes, they are related to independent sums of uniform distributions [BS22, Thm. 1.8]. Our aim in this section is to give another “generic” asymptotic normality criterion for sequences of CGF’s, which is sufficient to quickly identify the limit in many cases of interest.
Throughout the rest of the section, we will use the following notation. Let $a^{(N)}$ and $b^{(N)}$ for $N=1,2,\dots$ denote two sequences of multisets of positive integers of the same size. By 3.9(ii), we can always assume the values in $a^{(N)}$ are strictly greater than 1. The multisets $b^{(N)}$ may contain 1’s in order to have the same size as $a^{(N)}$. Let $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}/\prod_{b\in b^{(N)}}[b]_{q}$, and let $\mathcal{X}_{N}$ be the random variable associated to $\mathbb{E}[q^{\mathcal{X}_{N}}]=f_{N}(q)/f_{N}(1)$. We will denote the $d$th moment and cumulant of $\mathcal{X}_{N}$ by $\mu_{d}^{(N)}$ and $\kappa_{d}^{(N)}$ for $N=1,2,\ldots$ respectively. Similarly, the standard deviation and mean of $\mathcal{X}_{N}$ are denoted by $\sigma^{(N)}$ and $\mu^{(N)}$. Recall the notation for uniform sum distributions $\mathcal{S}_{\mathbf{t}}$ and $p$-norms from Section 2.4.
Theorem 4.1.
Let $a^{(N)}$ be a sequence of multisets of positive integers where $a\geq 2$ for each $a\in a^{(N)}$. Let $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}$ and $\mathcal{X}_{N}$ be the associated sequence of CGF polynomials and random variables.
@@ 4. Asymptotic considerations
    (4.1) $\sum_{a\in a^{(N)}}\left(\frac{a}{\max a^{(N)}}\right)^{4}\to\infty.$
@@
    Suppose the cardinality of the multisets $a^{(N)}$ is bounded by some $m$ and $\max a^{(N)}\to\infty$. Then $\mathcal{X}_{1}^{*},\mathcal{X}_{2}^{*},\ldots$ converges in distribution if and only if the rescaled multisets $a^{(N)}/|a^{(N)}|_{2}\in\widetilde{\ell}_{\leq m}$ converge pointwise to some multiset $\mathbf{t}\in\widetilde{\ell}_{\leq m}$ in the sense of [BS22, §3.2]. In that case, the limiting distribution is the standardized uniform sum $\mathcal{S}_{\mathbf{t}}^{*}$.
@@
Express $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}=\prod_{a\in a^{(N)}}[a]_{q}/[1]_{q}$ in rational form to obtain the corresponding sequence of multisets $b^{(N)}$ from the denominators. By 2.1 and the scaling property of cumulants, the standardized cumulants of $\mathcal{X}_{N}$ are given by
(4.2) $(\kappa_{d}^{(N)})^{*}=\frac{\kappa_{d}^{(N)}}{(\sigma^{(N)})^{d/2}}=\frac{\frac{B_{d}}{d}\sum_{a\in a^{(N)}}(a^{d}-1)}{\left(\frac{B_{2}}{2}\sum_{a\in a^{(N)}}(a^{2}-1)\right)^{d/2}}.$
Since $a\geq 2$ for all $a\in a^{(N)}$, $\frac{1}{2}a^{d}\leq a^{d}-1\leq a^{d}$, so $a^{d}-1=\Theta(a^{d})$ uniformly. Hence for $d\geq 2$ even,
@@
Therefore by 2.7, asymptotic normality is equivalent to $|a^{(N)}|_{4}/|a^{(N)}|_{2}\to 0$, or equivalently
@@
$\frac{|a^{(N)}|_{4}}{|a^{(N)}|_{\infty}}\leq\frac{|a^{(N)}|_{2}}{|a^{(N)}|_{4}}.$
Thus, $|a^{(N)}|_{4}/|a^{(N)}|_{\infty}\to\infty$ implies asymptotic normality. By Definition 2.10
@@
For (ii), let $\hat{a}^{(N)}\coloneqq\sqrt{12}\cdot a^{(N)}/|a^{(N)}|_{2}$ be the rescaled multiset such that $\mathcal{S}_{\hat{a}^{(N)}}$ is standardized for each $N$ according to (2.17). Since $|a^{(N)}|_{2}\geq|a^{(N)}|_{\infty}:=\max a^{(N)}\to\infty$, we have by (4.2) that
@@
Here and elsewhere we write $f(N)\sim g(N)$ to mean $\lim_{N\to\infty}f(N)/g(N)=1$. This last expression is the $d$th cumulant of $\mathcal{S}_{\hat{a}^{(N)}}$ by (2.17). Therefore, by the cumulant version of 2.5 and 5.2, we know $\mathcal{X}_{1}^{*},\mathcal{X}_{2}^{*},\ldots$ converges in distribution if and only if the sequence $\mathcal{S}_{\hat{a}^{(N)}}$ converges in distribution, and to the same limit. By [BS22, Thm. 1.15], $\mathcal{S}_{\hat{a}^{(N)}}$ converges in distribution if and only if the multisets $\hat{a}^{(N)}$ converge pointwise in the sense of [BS22, §3] to some $\mathbf{t}\in\tilde{\ell}_{2}$, with limiting distribution $\mathcal{S}_{\mathbf{t}}+\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{1-|\mathbf{t}|_{2}^{2}/12}$. Since the cardinality of $a^{(N)}$ is bounded, $|\hat{a}^{(N)}|_{2}^{2}=12$ for each $N$, so $|\mathbf{t}|_{2}^{2}=\lim_{N\to\infty}|a^{(N)}|_{2}^{2}=12$. This implies $\sigma=0$ by definition, so the limiting distribution of the standardized random variables $\mathcal{X}_{N}^{*}$ is just the standardized generalized uniform sum $\mathcal{S}_{\mathbf{t}}=\mathcal{S}_{\mathbf{t}}^{*}$ as given in the statement.
@@
Building on the case of finite products of $q$-integers above, we can now address the characterization of asymptotic normality for basic CGF’s stated in the introduction. Intuitively, the idea is that if the numerator of the rational form of the CGF leads to asymptotic normality, then the rational form does as well since it is dominated by its numerator.
Proof of 1.5.
Observe that for any positive integer $n$, the polynomial $[n]_{q}/n$ is the probability generating function of a discrete uniform random variable $\mathcal{U}_{n}$ supported on $\{0,1,\ldots,n-1\}$. Let $\mathcal{A}_{N}=\sum_{a\in a^{(N)}}\mathcal{U}_{a}$ and $\mathcal{B}_{N}=\sum_{b\in b^{(N)}}\mathcal{U}_{b}$ denote the CGF distributions corresponding with the numerator and denominator of $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}/\prod_{b\in b^{(N)}}[b]_{q}$. By 2.1, $\sigma_{\mathcal{A}_{N}}^{2}=\frac{1}{12}\sum_{a\in a^{(N)}}(a^{2}-1)$ and $\sigma_{\mathcal{B}_{N}}^{2}=\frac{1}{12}\sum_{b\in b^{(N)}}(b^{2}-1)$.
@@
(4.3) $\mathcal{X}_{N}+\mathcal{B}_{N}=\mathcal{A}_{N},$
@@
Let $c_{N}\coloneqq\sigma_{\mathcal{B}_{N}}/\sigma_{\mathcal{A}_{N}}$, so each $c_{N}\geq 0$. By the hypothesis in (1.5),
Recall that convergence in distribution of real-valued random variables can be metrized using the Lévy metric. Therefore, it suffices to show that every subsequence of $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ itself has an asymptotically normal subsequence. Hence, without loss of generality, we may assume that the sequence $c_{N}$ converges to some $10\leq c<$ 1.
@@
(4.4) $\mathcal{X}_{N}^{*}\ \sqrt{1-c_{N}^{2}}+\mathcal{B}_{N}^{*}\ c_{N}=\mathcal{A}_{N}^{*}.$
@@
(4.5) $\phi_{\mathcal{X}_{N}^{*}}(t)=\frac{\phi_{\mathcal{A}_{N}^{*}}(t\,/\sqrt{1-c_{N}^{2}})}{\phi_{\mathcal{B}_{N}^{*}}(t\,c_{N}/\sqrt{1-c_{N}^{2}})}.$
Since the hypothesis in (1.6) implies (4.1) holds, 4.1 says the sequence $\mathcal{A}_{N}$ is asymptotically normal, so by Lévy’s continuity theorem,
@@
for all $t\in\mathbb{R}$. This convergence is in fact uniform on bounded subsets of $\mathbb{R}$ (see e.g. [Bil95, Exercise 26.15(b)]). Consider the cases for $10\leq c<$ 1.
@@
    Suppose $c>0$. By 3.9(ii), we have $|b^{(N)}|_{\infty}\leq|a^{(N)}|_{\infty}$. This observation and the fact that $c>0$ imply that (1.6) holds with $a^{(N)}$ replaced by $b^{(N)}$. Hence
@@
    the result follows from (4.5) and the calculations from the previous case.
Example 4.2.
@@
Let $x=k/n$. Since both $k,n-k\to\infty$ as $N\to\infty$, we may suppose $k\leq n-k$ for all $N$ since we can replace $k$ with $n-k$ if necessary without changing $f_{N}(q)$, so $20\leq x\leq 1/$ 2. It is easy to check that $x^{3}/(1-(1-x)^{3})\leq 1/7$ for $20\leq x\leq 1/$ 2, so that (1.5) holds. We also see
@@
so (1.6) holds as $k\to\infty$ since $(5-10x+10x^{2}-5x^{3}+x^{4})$ is positive for all real values in the range $20\leq x\leq 1/$ 2. Thus, 1.5 applies to show the sequence of CGF random variables corresponding to $f_{N}(q)$ is asymptotically normal provided $k,n-k\to\infty$. The details of this calculation are left implicit in [Dia88].
5. Analytic considerations
In this section, we consider the problem of classifying all standardized CGF distribution via analytic considerations of their characteristic functions $\phi_{\mathcal{X}^{*}}(t)$ and their second characteristic functions, $\log\phi_{\mathcal{X}^{*}}(t)$. Given the many different characterizations of CGF’s introduced in 1.1, we have a rich set of tools for studying these complex functions. In particular, we will show that the limiting standardized characteristic functions are entire, see 5.3. This allows us to complete the proof of 1.8. We begin by spelling out the connections between the CGF’s in rational and cyclotomic form with the first and second (standardized) characteristic functions.

### v3

@@ 3.5 CGFs with a fixed denominator
Lemma 35.
Fix a multiset $B=\{b_{1},\ldots,b_{m}\}$ of positive integers. Suppose $\{a_{1},\ldots,a_{m}\}$ and $\{a_{1}^{\prime},\ldots,a_{m}^{\prime}\}$ are multisets of positive integers such that both $\prod_{k=1}^{m}[a_{k}]_{q}/[b_{k}]_{q}$ and $\prod_{k=1}^{m}[a_{k}^{\prime}]_{q}/[b_{k}]_{q}\in\Phi^{+}$ are cyclotomic generating functions. Then there exists a multiset $\{h_{1},\ldots,h_{m}\}$ of positive integers such that for all $1\leqslant i\leqslant m$,
$\frac{\prod_{k=1}^{i}[h_{k}]_{q}\prod_{k=i+1}^{m}[a_{k}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}\in\Phi^{+}\qquad\text{and}\qquad\frac{\prod_{k=1}^{i}[h_{k}]_{q}\prod_{k=i+1}^{m}[a_{k}^{\prime}]_{q}}{\prod_{k=1}^{m}[b_{k}]_{q}}\in\Phi^{+}.$
@@
By long division, one may observe $[xy]_{q}/[y]_{q}=[x]_{q^{y}}\in\Phi^{+}$. So, we may replace any $[a_{k}]_{q}$ in the numerator of a cyclotomic generating function with $[\ell a_{k}]_{q}$ for any positive integer $\ell$ and get another cyclotomic generating function. Hence, defining $h_{k}=a_{k}a_{k}^{\prime}$ for $1\leqslant k\leqslant m$ has the desired property. ∎
Corollary 36.
@@
4 Asymptotic considerations
Given a sequence of cyclotomic generating functions and their corresponding random variables, we can put their distributions in one common frame of reference by standardizing them to have mean 0 and standard deviation 1. Under what conditions does such a sequence converge? Can we classify all possible standardized limiting distributions of random variables coming from CGFs? Both problems have been completely solved for certain families of CGFs coming from $q$-hook length formulas, but the complete classification is not known, see [BS22, Open Problem 1.19]. In several “generic” regimes, they are asymptotically normal [BKS20a, Thm. 1.7], [BS22, Thm. 1.13]. In mildly degenerate regimes, they are related to independent sums of uniform distributions [BS22, Thm. 1.8]. Our aim in this section is to give another “generic” asymptotic normality criterion for sequences of CGFs, which is sufficient to quickly identify the limit in many cases of interest.
Throughout the rest of the section, we will use the following notation. Let $a^{(N)}$ and $b^{(N)}$ for $N=1,2,\dots$ denote two sequences of multisets of positive integers of the same size. By Lemma 30(ii), we can always assume the values in $a^{(N)}$ are strictly greater than 1. The multisets $b^{(N)}$ may contain 1’s in order to have the same size as $a^{(N)}$. Let $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}/\prod_{b\in b^{(N)}}[b]_{q}$, and let $\mathcal{X}_{N}$ be the random variable associated to $\mathbb{E}[q^{\mathcal{X}_{N}}]=f_{N}(q)/f_{N}(1)$. We will denote the $d$th moment and cumulant of $\mathcal{X}_{N}$ by $\mu_{d}^{(N)}$ and $\kappa_{d}^{(N)}$ for $N=1,2,\ldots$ respectively. Similarly, the standard deviation and mean of $\mathcal{X}_{N}$ are denoted by $\sigma^{(N)}$ and $\mu^{(N)}$. Recall the notation for uniform sum distributions $\mathcal{S}_{\mathbf{t}}$ and $p$-norms from Section 2.4.
Theorem 37.
Let $a^{(N)}$ be a sequence of multisets of positive integers where $a\geqslant 2$ for each $a\in a^{(N)}$. Let $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}$ and $\mathcal{X}_{N}$ be the associated sequence of CGF polynomials and random variables.
@@ 4 Asymptotic considerations
    $\sum_{a\in a^{(N)}}\left(\frac{a}{\max a^{(N)}}\right)^{4}\to\infty.$ (4.1)
@@
    Suppose the cardinality of the multisets $a^{(N)}$ is bounded by some $m$ and $\max a^{(N)}\to\infty$. Then $\mathcal{X}_{1}^{*},\mathcal{X}_{2}^{*},\ldots$ converges in distribution if and only if the rescaled multisets $a^{(N)}/|a^{(N)}|_{2}\in\widetilde{\ell}_{\leqslant m}$ converge pointwise to some multiset $\mathbf{t}\in\widetilde{\ell}_{\leqslant m}$ in the sense of [BS22, §3.2]. In that case, the limiting distribution is the standardized uniform sum $\mathcal{S}_{\mathbf{t}}^{*}$.
@@
The polynomial $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}$ can also be expressed in rational form as $=\prod_{a\in a^{(N)}}[a]_{q}/[1]_{q}$. Thus, we obtain the corresponding sequence of multisets $b^{(N)}=(1,1,\dots,1)$ from the denominators where the size of $b^{(N)}$ equals the size of $a^{(N)}$ as multisets. By Theorem 9 and the scaling property of cumulants, the standardized cumulants of $\mathcal{X}_{N}$ are given by
$(\kappa_{d}^{(N)})^{*}=\frac{\kappa_{d}^{(N)}}{(\sigma^{(N)})^{d/2}}=\frac{\frac{B_{d}}{d}\sum_{a\in a^{(N)}}(a^{d}-1)}{\left(\frac{B_{2}}{2}\sum_{a\in a^{(N)}}(a^{2}-1)\right)^{d/2}}.$ (4.2)
Since $a\geqslant 2$ for all $a\in a^{(N)}$, $\frac{1}{2}a^{d}\leqslant a^{d}-1\leqslant a^{d}$, so $a^{d}-1=\Theta(a^{d})$ uniformly. Hence for $d\geqslant 2$ even,
@@
Therefore by Theorem 15, asymptotic normality is equivalent to $|a^{(N)}|_{4}/|a^{(N)}|_{2}\to 0$, or equivalently
@@
$\frac{|a^{(N)}|_{4}}{|a^{(N)}|_{\infty}}\leqslant\frac{|a^{(N)}|_{2}}{|a^{(N)}|_{4}}.$
Thus, $|a^{(N)}|_{4}/|a^{(N)}|_{\infty}\to\infty$ implies asymptotic normality. By Definition 18
@@
For (ii), let $\hat{a}^{(N)}\coloneqq\sqrt{12}\cdot a^{(N)}/|a^{(N)}|_{2}$ be the rescaled multiset such that $\mathcal{S}_{\hat{a}^{(N)}}$ is standardized for each $N$ according to (2.17). Since $|a^{(N)}|_{2}\geqslant|a^{(N)}|_{\infty}:=\max a^{(N)}\to\infty$, we have by (4.2) that
@@
Here and elsewhere we write $f(N)\sim g(N)$ to mean $\lim_{N\to\infty}f(N)/g(N)=1$. This last expression is the $d$th cumulant of $\mathcal{S}_{\hat{a}^{(N)}}$ by (2.17). Therefore, by the cumulant version of Theorem 13 and Theorem 40, we know $\mathcal{X}_{1}^{*},\mathcal{X}_{2}^{*},\ldots$ converges in distribution if and only if the sequence $\mathcal{S}_{\hat{a}^{(N)}}$ converges in distribution, and to the same limit. By [BS22, Thm. 1.15], $\mathcal{S}_{\hat{a}^{(N)}}$ converges in distribution if and only if the multisets $\hat{a}^{(N)}$ converge pointwise in the sense of [BS22, §3] to some $\mathbf{t}\in\tilde{\ell}_{2}$, with limiting distribution $\mathcal{S}_{\mathbf{t}}+\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{1-|\mathbf{t}|_{2}^{2}/12}$. Since the cardinality of $a^{(N)}$ is bounded, $|\hat{a}^{(N)}|_{2}^{2}=12$ for each $N$, so $|\mathbf{t}|_{2}^{2}=\lim_{N\to\infty}|a^{(N)}|_{2}^{2}=12$. This implies $\sigma=0$ by definition, so the limiting distribution of the standardized random variables $\mathcal{X}_{N}^{*}$ is just the standardized generalized uniform sum $\mathcal{S}_{\mathbf{t}}=\mathcal{S}_{\mathbf{t}}^{*}$ as given in the statement.
@@
Building on the case of finite products of $q$-integers above, we can now address the characterization of asymptotic normality for basic CGFs stated in the introduction. Intuitively, the idea is that if the numerator of the rational form of the CGF leads to asymptotic normality, then the rational form does as well since it is dominated by its numerator.
Proof of Theorem 5.
Observe that for any positive integer $n$, the polynomial $[n]_{q}/n$ is the probability generating function of a discrete uniform random variable $\mathcal{U}_{n}$ supported on $\{0,1,\ldots,n-1\}$. Let $\mathcal{A}_{N}=\sum_{a\in a^{(N)}}\mathcal{U}_{a}$ and $\mathcal{B}_{N}=\sum_{b\in b^{(N)}}\mathcal{U}_{b}$ denote the CGF distributions corresponding to the numerator and denominator of $f_{N}(q)=\prod_{a\in a^{(N)}}[a]_{q}/\prod_{b\in b^{(N)}}[b]_{q}$. By Theorem 9, $\sigma_{\mathcal{A}_{N}}^{2}=\frac{1}{12}\sum_{a\in a^{(N)}}(a^{2}-1)$ and $\sigma_{\mathcal{B}_{N}}^{2}=\frac{1}{12}\sum_{b\in b^{(N)}}(b^{2}-1)$.
Let $c_{N}\coloneqq\sigma_{\mathcal{B}_{N}}/\sigma_{\mathcal{A}_{N}}$, so each $c_{N}\geqslant 0$. By the hypothesis in (1.5),
Recall that convergence in distribution of real-valued random variables can be metrized using the Lévy metric. Therefore, it suffices to show that every subsequence of $\mathcal{X}_{1},\mathcal{X}_{2},\ldots$ itself has an asymptotically normal subsequence. Hence, without loss of generality, we may assume that the sequence $c_{N}$ converges to some $10\leqslant c<$ 1.
@@
$\mathcal{X}_{N}+\mathcal{B}_{N}=\mathcal{A}_{N},$ (4.3)
@@
$\mathcal{X}_{N}^{*}\ \sqrt{1-c_{N}^{2}}+\mathcal{B}_{N}^{*}\ c_{N}=\mathcal{A}_{N}^{*}.$ (4.4)
@@
$\phi_{\mathcal{X}_{N}^{*}}(t)=\frac{\phi_{\mathcal{A}_{N}^{*}}\left(t\,/\sqrt{1-c_{N}^{2}}\right)}{\phi_{\mathcal{B}_{N}^{*}}\left(t\,c_{N}/\sqrt{1-c_{N}^{2}}\right)}.$ (4.5)
Since the hypothesis in (1.6) implies (4.1) holds, Theorem 37 says the sequence $\mathcal{A}_{N}$ is asymptotically normal, so by Lévy’s continuity Theorem 12,
@@
for all $t\in\mathbb{R}$. This convergence is in fact uniform on bounded subsets of $\mathbb{R}$ (see e.g. [Bil95, Exercise 26.15(b)]). Consider the cases for $10\leqslant c<$ 1.
@@
    Suppose $c>0$. By Lemma 30(ii), we have $|b^{(N)}|_{\infty}\leqslant|a^{(N)}|_{\infty}$. This observation and the fact that $c>0$ imply that (1.6) holds with $a^{(N)}$ replaced by $b^{(N)}$. Hence
@@
    and the result follows from (4.5) and the calculations from the previous case.∎
Example 38.
@@
Let $x=k/n$. Since both $k,n-k\to\infty$ as $N\to\infty$, we may suppose $k\leqslant n-k$ for all $N$ since we can replace $k$ with $n-k$ if necessary without changing $f_{N}(q)$, so $20\leqslant x\leqslant 1/$ 2. It is easy to check that $x^{3}/(1-(1-x)^{3})\leqslant 1/7$ for $20\leqslant x\leqslant 1/$ 2, so that (1.5) holds. We also see
@@
so (1.6) holds as $k\to\infty$ since $(5-10x+10x^{2}-5x^{3}+x^{4})$ is positive for all real values in the range $20\leqslant x\leqslant 1/$ 2. Thus, Theorem 5 applies to show the sequence of CGF random variables corresponding to $f_{N}(q)$ is asymptotically normal provided $k,n-k\to\infty$. The details of this calculation are left implicit in [Dia88].
5 Analytic considerations
In this section, we consider the problem of classifying all standardized CGF distribution via analytic considerations of their characteristic functions $\phi_{\mathcal{X}^{*}}(t)$ and their second characteristic functions, $\log\phi_{\mathcal{X}^{*}}(t)$. Given the many different characterizations of CGFs introduced in 1, we have a rich set of tools for studying these complex functions. In particular, we will show that the limiting standardized characteristic functions are entire, see Corollary 41. This allows us to complete the proof of Theorem 8. We begin by spelling out the connections between the CGFs in rational and cyclotomic form with the first and second (standardized) characteristic functions.